### v1

# Investigating Large Language Models in Diagnosing Students’ Cognitive Skills in Math Problem-solving

Abstract
[…]
1 Introduction
Automatic assessment of students’ responses to mathematical problems can support learning at scale. Assessment can serve as feedback to guide students through problem-solving processes, allowing them to refine their reasoning and correct misconceptions as they arise. As teachers’ manual assessment is limited in scale and availability (Stankous, 2016), automated assessment is particularly instrumental in environments where human instructors are unavailable or where student-to-teacher ratios are too high (e.g., MOOCs).
The advancement in AI technologies has expanded the automatic assessment to evaluate more complicated forms of student responses. While early automated assessment methods worked on grading categorical multiple-choice questions, current multimodal and reasoning AI models (e.g., GPT-o1) can process handwritten responses to constructed response questions (Gao et al., 2024; Liu et al., 2024; Chamieh et al., 2024; Baral et al., 2023). As opposed to multiple choices and short answers, these students’ handwritten constructed responses capture students’ thinking processes (Livingston, 2009).
Evaluating students’ constructed responses with AI presents a promising opportunity to diagnose their cognitive skills at scale. Cognitive skills are the abilities to know, apply, and reason mathematical concepts and formulas (Mullis, 2017). Unlike traditional grading that focuses primarily on final answers, cognitive diagnosis enables more granular feedback by identifying where a student’s thinking process may have broken down (Jin et al., 2024; Rahbarnia et al., 2014; Kheong, 1994). For instance, such evaluation can reveal whether a student successfully recalled relevant concepts, represented a problem with an appropriate model, determined a reasonable strategy, and implemented it correctly. These cognitive skills (Mullis, 2017) (complete list in Table 3) often manifest in varied ways through texts and graphics and in subtle ways without overt evidence in students’ responses.
However, such an in-depth diagnosis is challenging because AI models need to understand complicated handwritten responses and need to infer students’ thinking processes beyond visible cues. This diagnosis task involves perceiving graphics in students’ responses, interpreting mathematical notations, and reasoning about students’ problem-solving processes. Existing work has focused on examining AI models’ performance on each of these capabilities in math problem-solving (Didolkar et al., 2024; Fang et al., 2024; Ahn et al., 2024), multimodal perception (Zhang et al., 2024b; a), and reasoning (Wang et al., 2023; Hao et al., 2024) in diverse domains. Whereas previous research has asked how well LLMs can emulate human cognition, our task inverts that perspective, challenging models to infer humans’ logical cognition from partial visual and textual evidence. The diagnosis task is precisely the intersection of these capabilities that makes it particularly complex for current LLMs.
In this work, we investigate how well existing LLMs diagnose students’ cognitive skills in mathematics. Specifically, we address two research questions:
  1. RQ1.
    How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
  2. RQ2.
    How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
To answer these questions, we constructed MathCog, a benchmark dataset designed to evaluate cognitive skill diagnosis. In collaboration with 5 education experts and 15 middle school teachers, we curated 12 middle school math problems, each with 50 student responses diagnosed by teachers based on a diagnostic checklist. Each checklist is based on a well-founded diagnosis framework (Mullis, 2017) and consists of 4 to 6 cognitive skill check items (see Table 1).
[…]
  • Creating a benchmark dataset for evaluating cognitive skill diagnosis performance, MathCog, composed of expert-reviewed diagnostic checklists and teacher-diagnosed results for 639 student responses across 12 middle school math problems.
  • Examining the performance of 16 state-of-the-art closed and open LLMs on diagnosing students’ cognitive skills using MathCog.
  • Investigating the impact of multimodal input, reasoning, and model size on the cognitive skill diagnosis task through a comparison study.
2 Related Work
We review the foundational framework underlying our cognitive diagnosis task, followed by prior investigations into LLMs’ performance in math-related contexts, and existing benchmarks, highlighting both their relevance and current gaps.
2.1 Cognitive Diagnosis in Mathematics Education
The TIMSS (Trends in International Mathematics and Science Study) framework (Mullis, 2017) provides a comprehensive and mathematics-specific approach to cognitive diagnosis, comprising content and cognitive domains. The cognitive domain, which evaluates knowledge application, is divided into three key areas: Knowing (recalling definitions, recognizing mathematical entities, computing operations), Applying (determining strategies, representing problems, implementing solutions), and Reasoning (analyzing, integrating, evaluating, justifying). While other frameworks like Bloom’s taxonomy (Bloom et al., 1956) and Polya’s problem-solving framework (Polya, 1957) exist, TIMSS’s mathematics-specific approach, validated through its extensive use across approximately 70 countries since 1995 (Mullis et al., 2023), provides more targeted and reliable guidance for assessing students’ mathematical thinking. Our cognitive skill diagnosis task is founded on this framework that provides deeper insights into students’ mathematical thinking processes, addresses limitations of knowledge-centered assessments, and offers specific guidelines for educational interventions.
2.2 LLM Capabilities in Mathematical Tasks
LLMs have shown significant progress in mathematical problem-solving, with advanced models achieving improved reasoning capabilities through reinforcement learning (Guo et al., 2025; OpenAI, 2024; Zhong et al., 2024). While larger models and Chain-of-Thought prompting further enhance mathematical reasoning capabilities (Li et al., 2024; Yang et al., 2024a; Wei et al., 2022), current models still struggle with multimodal inputs such as visual elements and handwritten content (Zhang et al., 2024b; Baral et al., 2025; Liu et al., 2024). Although domain-specific models like MathGLM-Vision (Yang et al., 2024b) show promise in addressing these challenges, their performance remains limited by data scarcity (Yan et al., 2024). Beyond direct problem-solving, LLMs need metacognitive abilities to evaluate student responses and identify misconceptions (Baral et al., 2024; Kaggle & Eedi, 2020), yet current research focuses mainly on basic error identification rather than complex cognitive diagnosis. We address a critical gap in the field by providing a comprehensive evaluation of LLMs’ ability to track thinking processes and assess cognitive abilities in mathematical contexts.
2.3 Benchmarks for Evaluating Mathematical Tasks
A diverse array of benchmarks exists for evaluating LLMs’ mathematical capabilities, including GSM8K (Cobbe et al., 2021) for elementary school problems, MATH (Hendrycks et al., 2021) for high school competition questions, MathOdyssey (Fang et al., 2024), and DeepMind’s Mathematics dataset (Saxton et al., 2019). In the educational domain, specialized assessments have emerged, such as MathFish (Lucy et al., 2024), which aligns 9,900 problems with 385 K-12 standards to evaluate models’ capacity to recognize specific skills and concepts, along with handwriting recognition (Baral et al., 2025) and misconception detection in MCQs (Kaggle & Eedi, 2020). However, with few exceptions like (Hellman et al., 2023), there is a notable gap in datasets designed for cognitive diagnosis in open-ended responses. While existing research has mainly focused on simple error identification and scoring, the diagnosis of students’ complex cognitive processes in open-ended questions has been relatively less explored. For our novel cognitive skill diagnosis task, we constructed a dataset that supports a comprehensive evaluation of LLMs’ ability to track implicit thinking processes, identify root causes of errors, and assess students’ cognitive abilities.
3 Dataset: MathCog
To evaluate LLMs’ performance on the cognitive skill diagnosis, we first created a benchmark dataset composed of secondary school math problems, student responses, diagnostic checklists, and verdicts (Table 1). The math problem and student response data are from AI-Hub [^1] https://www.aihub.or.kr, which contains 5,932 K-12 math problem types and 147,882 handwritten student response images labeled with OCR transcriptions. Each problem type has 5 to 7 isomorphic problems that have the same problem-solving procedures and differ only in the numbers in the problem description (e.g., “Solve $x^{2}+2x-3=0$” and “Solve $x^{2}+x-2=0$”). Among the data, we chose the problem types and student responses corresponding to grades 7-9, as the problems were complex enough to elicit cognitive processes. Among 2,605 problem types from grades 7 to 9, we excluded the problem types that had fewer than 50 student responses to ensure that we had enough student responses to use in our performance evaluation. Finally, we manually reviewed the remaining 35 problem types and removed the problem types whose corresponding isomorphic problems require different mathematical knowledge (e.g., “Solve $x^{2}+2x-3=0$” (quadratic) vs. “Solve $x^{3}+4x-1=0$” (trinomial)). The filtering process left us with 137 problems and 796 student responses from 15 problem types, each involving the same mathematical knowledge and problem-solving procedures. The specific numbers at each stage of filtering are shown in Figure 1.
Figure 1: The dataset creation process and the number of problem types, problems, and student responses at each filtering stage.
Problem Student Response Diagnostic Checklist Verdict
When $P=\sqrt{3}$ $-7\sqrt{5}+2\sqrt{3}$, $Q=2\sqrt{3}-\sqrt{5}$ $+2\sqrt{5}$, find the value of $P+Q$. [Uncaptioned image]
Recognize: Does the student
realize that if the numbers in
the radicals are the same, he
or she can perform the four
arithmetic operations on the
coefficients with the same
letters?
Vague
Yes
Compute: Were the addition
and subtraction calculations
of coefficients performed
accurately?
Evident
Yes
Determine: Does the student
choose a strategy to first
simply organize each equation
(e.g., put $\sqrt{5}$ terms together,
etc.) and then find the final
sum?
Evident
No
There is a trapezoid with a lower side length of $8$ $cm$ and a height of $4$ $cm$. If the area of this trapezoid is not less than $28$ $cm^{2}$, find how much more $cm$ the length of the upper side of the trapezoid must be. [Uncaptioned image]
Recall: Does the student
remember the formula for
the area of a shape correctly
(e.g., the formula for the area
of a trapezoid)?
Evident
Yes
Compute: Has the student
calculated the linear and
constant terms correctly?
Vague
No
Determine: Did the student
know the need to set up an
equation and then solve it
to find the range of solutions
that meet the conditions?
Evident
Yes
Represent: Has the given
situation been expressed
correctly?
Evident
No
Implement: When simplifying
an expression, does the student
keep the expression correct by
performing the same operation
on both sides?
Vague
Yes
Table 1: Two samples from MathCog. Each data point is composed of a math problem, student response, relevant diagnostic checklist, and verdict for each check item.
3.1 Diagnostic Checklist
For each problem type, we made a diagnostic checklist commonly applicable to isomorphic problems. These checklists are a list of binary question items mapped to one of the 15 cognitive skills defined in the TIMSS 2019 assessment framework (Mullis, 2017). The authors initially drafted the checklist items and refined them by requesting a review from five math curriculum and evaluation experts with PhD degrees in education and practical experience in making math assessment guidelines (e.g., TIMSS, Korean public school curriculum). Experts gave feedback on the clarity, granularity, and validity of the checklist items. For example, the experts suggested changing “square root term” into “irrational term” for generality and combining items like “Were coefficients calculated correctly?” and “Were addition and subtraction performed correctly?” into “Were the addition and subtraction calculations of coefficients performed accurately?” for reasonable granularity: specific enough to describe a distinct skill, yet general enough to accommodate diverse forms of valid evidence. The experts also commented on the skills each problem can or cannot assess. Experts pointed out that our math problems primarily focus on calculating numbers and applying knowledge and, hence, are limited in assessing “reasoning” (e.g., justifying, analyzing, generalizing) (Mullis, 2017) by their design. We scoped our check items to “knowing” and “applying” cognitive domains only. We took two iterations to refine the checklists, and each checklist was reviewed by two experts independently in each iteration.
3.2 Teacher-generated Verdict
We recruited 15 middle school math teachers to evaluate 796 student responses based on predefined diagnostic checklists. These teachers had an average of 6.1±4.3 years of experience (range: 2.5–20 years). Each check item was assessed along two dimensions: “yes/no” and “evident/vague” (see Table 1). A “yes” response indicated that the student fully demonstrated the cognitive action specified, while a “no” indicated otherwise. “Evident” meant there was clear evidence to support the judgment, whereas “vague” signified insufficient evidence. For example, for the check item “Did the student calculate correctly?”, teachers marked “evident yes” if the response showed a complete, error-free calculation process, “evident no” if there were clear mistakes, and “vague no” if calculations were missing.
To account for the subjectivity in the diagnosis, we grouped three teachers to evaluate approximately 160 student responses together. The diagnosis occurred in two phases: first, each group collaboratively assessed 36 responses, resolving conflicts through majority voting; second, each teacher independently assessed 18 responses. The second phase included 18 overlapping responses, allowing us to measure inter-rater agreement (see Table  4 in Appendix). Based on the threshold established in prior literature (Graham et al., 2012), we set a minimum agreement of 70% and excluded data from problem types 6, 13, and 14 accordingly. The distribution of the verdicts is presented in Table 5.
4 Experimental Setting
Using MathCog as a benchmark dataset, we explored two research questions. For RQ1, we evaluated the state-of-the-art LLMs, testing GPT (Achiam et al., 2023), Gemini (Team et al., 2023), Llama (Touvron et al., 2023), Claude (Anthropic, 2024), and DeepSeek (Guo et al., 2025) series. For RQ2, we configured three experiments in which we compared the performance of LLMs in different inputs, reasoning capabilities, and model sizes.
Task Definition. Our task for LLMs is essentially a single-label classification task. Formally, we represent diagnosis result for a student’s response as a collection of tuples $(c,r)$, where $c\in C$ (i.e., check items) and $r\in\{Evident\ Yes,Vague\ Yes,Evident\ No,Vague\ No\}$. The combined judgment $(c,r)$ provides both an evaluative decision and a confidence signal for a skill. Given a context $(P,S,R,C)$, where P is the problem, S is the answer, and R is the student response, the task is to produce assessment tuples $(c,r)$ for all $c\in C$.
Prompting. We instructed LLMs to evaluate each check item based on a given math problem, its solution, a student’s response, and a diagnostic checklist. The LLMs processed inputs in the form of OCR transcriptions, where mathematical formulas and visual cues (e.g., strikethroughs) were represented using LaTeX. In the case of testing the effect of multimodality, we supplied images of student responses. Since the original inputs were in Korean, we machine-translated them into English to prevent possible performance degradation due to language (Achiam et al., 2023). To enhance response accuracy, we employed Chain-of-Thought prompting (Wei et al., 2022; Liu et al., 2024), guiding the LLM to systematically address each check item by first restating its content, identifying relevant evidence, providing an explanation, and delivering a final verdict. The verdict followed one of the four categories used by teachers in MathCog. We opted not to include few-shot examples, as MathCog does not provide explicit evidence and explanations for reference. To minimize randomness, all LLMs were run with a temperature setting of zero. A detailed description of the system and user prompts can be found in the Appendix.
Models. We conducted our experiment using 16 distinct closed and open LLMs from various vendors. To assess the impact of multimodal input, we tested GPT-4o (2024-08-06), Claude-3.5-Sonnet (20240620), and Gemini-1.5-Flash under two conditions: one using text-only prompts and another supplemented with an image. To examine the effect of reasoning capabilities, we compared LLMs with reasoning capabilities—GPT-o1-Preview (2024-09-12), Gemini-2.0-Flash-Thinking (exp-01-21), and DeepSeek-R1—with conventional LLMs of the same model series—GPT-4o, GPT-4o-Mini (2024-07-18), Gemini-1.5-Pro (002), Gemini-1.5-Flash, and DeepSeek-V3 (0324). Finally, we investigated model size differences: GPT-4o, Gemini-1.5-Pro, Llama-3.1-450B (2024-07-23), and DeepSeek-V3 were categorized as large models; Llama-3.3-70B (2024-07-23) and Qwen-2.5-72B (2024-09) as medium models; and GPT-4o-Mini, Gemini-1.5-Flash-8B (001), Llama-3.1-8B (2024-07-23), Qwen-2.5-7B (2024-09), and Mistral-7B (2024-08-22) as small models.
Metrics. We evaluated the outputs of the large language models (LLMs) against ground-truth diagnostic results provided by human experts. Performance was assessed using both macro F1 score and accuracy computed over verdicts (i.e., $(c,r)$) to quantify the effectiveness of skill diagnosis. For a more nuanced analysis, we also examined the models’ tendencies toward overconfidence and underconfidence. Specifically, overconfidence is defined as the conditional proportion of “evident” verdicts produced by the LLMs that were incorrect. Conversely, underconfidence is the conditional proportion of verdicts labeled as “vague” by the LLMs that were, in fact, both correct and should have been classified as “evident.”
5 Results
In this section, we organize and interpret the findings to answer our research questions. The complete results from our experiment are presented in Table 6 in the Appendix.
5.1 RQ1. How accurately do state-of-the-art LLMs diagnose students’ cognitive skills?
Overall, the results underscore the difficulty of the task (Table 2); no model surpassed an F1 score of 0.5, indicating that even state-of-the-art models struggle to diagnose students’ cognitive skills accurately. Our manual error analysis on a few samples revealed that LLMs largely failed when 1) student response flows in a non-directional, scattered layout, 2) a student uses underscores, strikethroughs, and calculation notes, and 3) a response lacks direct evidence for a skill, despite contextual information being sufficient for inference.
[…]
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-V3 .406 .461 .417 .714 .818 .107
GPT-4o .398 .476 .412 .672 .788 .110
Llama-3.1-405B .415 .438 .384 .624 .778 .186
Claude-3.5-Sonnet .329 .387 .332 .654 .812 .173
Gemini-1.5-Pro .430 .496 .441 .707 .840 .132
Table 2: The performance of the five state-of-the-art LLMs on the cognitive skill diagnosis task.
Analysis of skill-specific performance (Figure 2) reveals further insights into the limitations of current LLMs. No skill category achieved a maximum F1 score above 0.5. Interestingly, the skill Compute, which most closely resembles conventional grading tasks (i.e., evaluating the correctness of mathematical procedures and calculations), still yielded only a moderate F1 score across models. This moderate F1 score confirms LLMs’ struggle in verifying correct computation in constructed student responses.
Contrary to expectations, there was no substantial performance gap between the “Knowing” and “Applying” cognitive skills. Although we anticipated that “Knowing” skills (e.g., Recall, Recognize) would be easier than “Applying” skills for LLMs due to their surface-level nature, models performed comparably. An exception is found in the skill Classify/Order, which shows a relatively high average accuracy but a lower F1 score. Closer inspection reveals that “Evident Yes” is dominant in the ground truth dataset (Table 5), leading to the discrepancy between F1 score and accuracy.
Figure 2: The performance of five state-of-the-art LLMs across cognitive skill types. The left plot shows the minimum (bottom $\times$), maximum (top $\times$), and average (middle •) F1 scores for each skill category, while the right plot presents the corresponding accuracy metrics.
5.2 RQ2. How do multimodal input, reasoning, and model size influence LLMs’ cognitive skill diagnosis?
We observed subtle impacts of multimodal input and reasoning on the performance, while the impact of model size was significant.
Multimodal Input. Models supporting image input (e.g., GPT-4o-img, Claude-3.5-Sonnet-img, Gemini-1.5-Flash-img) consistently outperformed their text-only counterparts in accuracy (up to 10.6%) and in reducing underconfidence (up to 4.9%) (Figure 3 left), although the statistical difference between the pairs was not significant. Image input was particularly beneficial in cases where OCR transcription introduced errors. For example, a student underlined an arithmetic expression and wrote the corresponding calculation beneath it (see Figure 5 a). The OCR system misinterpreted this layout as a fraction, leading Claude-3.5-Sonnet (text-only) to incorrectly conclude that the student had set up an invalid equation. In contrast, the same model with image input correctly inferred the student’s intention and produced an accurate diagnosis.
Reasoning. The performance of reasoning-oriented models showed mixed results (Figure 3 right). DeepSeek-R1 achieved the highest F1 and accuracy scores among all models. However, other reasoning models, such as o1-Preview and Gemini-2.0-Flash-Thinking, did not show clear advantages over standard models in either accuracy or F1 score. These results suggest that reasoning capability alone does not guarantee better performance. For instance (Figure 5 b), DeepSeek-R1 classified a response as “Vague No” because the student omitted the factorization steps. However, the original expression was $\sqrt{48a}$, and the model failed to detect a clear error in the variable term, indicating that the student’s reasoning was evidently flawed. Throughout its reasoning process, the model focused on the constant term only and overlooked other possible places of error. While current reasoning models excel at pushing forward linear, step-by-step logic, effective cognitive diagnosis may also require divergent, multi-perspective reasoning for thorough analysis.
[…]
Figure 3: Impact of multimodal input (left) and reasoning capability (right) on LLMs’ performance in cognitive skill diagnosis. On the left, each adjacent pair of bars shows performance with and without the image input from the same model family. On the right, models with reasoning capabilities are grouped at the front.
Figure 4: Impact of model size on LLM performance in cognitive skill diagnosis. Models are grouped into large, medium, and small categories.
6 Discussion
Our findings imply that current LLMs are not yet suitable for high-stakes cognitive skill diagnosis. The models exhibit low recall and high overconfidence (Table 2), raising concerns that LLM-based systems may frequently overlook students’ misconceptions. This shortcoming is particularly problematic given that the primary goal of automated assessment is to identify and address such misconceptions, and it can compromise teachers’ and students’ trust in automated systems (Shin, 2021). To responsibly integrate LLMs into real-world assessment settings, practitioners should begin by evaluating their performance on small, ground-truth samples. This process not only exposes limitations but also allows for fine-tuning model judgments and confidence levels for downstream math topics and student populations. In parallel, model developers may work to mitigate overconfidence during training by encouraging more conservative, evidence-based decision-making.
Our manual analysis of failed cases reveals that inferring students’ thought processes from handwritten responses presents distinct challenges. While mathematical problem-solving is inherently procedural, students often express their reasoning in non-linear, spatially scattered layouts that lack clear directionality. This ill-structured presentation differs significantly from the format of most existing reasoning and problem-solving benchmarks. Moreover, ambiguous notations (e.g., fractions misinterpreted as underscores) and unconventional layouts frequently result in OCR transcription errors, which hinder accurate diagnosis unless image inputs are provided. To enable more reliable inference, future work may require richer representations of math responses that capture high-level semantic intent beyond surface-level positional and textual cues.
7 Limitations and Future Work
While MathCogrepresents a novel benchmark for evaluating LLMs on cognitive skill diagnosis, it has several limitations. First, due to data-sharing restrictions from the source provider (AIHub), the original math problems and student responses cannot be publicly released outside Korea. To support future research despite this constraint, we release the complete set of diagnostic checklists and expert-generated human annotations, which can serve as references for researchers to develop similar datasets in their educational contexts. In parallel, we are actively collecting new student response data under informed consent, with the goal of constructing a fully public benchmark in the future.
This ongoing data collection effort will also address limitations in the diversity of problem types represented in MathCog. Currently, the dataset focuses on nine problem types, primarily centered around arithmetic equation solving. Other areas of mathematics—such as geometry—often require interpreting visual representations and may further benefit from the integration of multimodal input. Additionally, reasoning-oriented problems, which explicitly ask students to generalize mathematical knowledge or justify their arguments, are underrepresented. We plan to expand the dataset to include a broader range of content areas and grade levels, enabling a more holistic evaluation of LLM performance across the mathematics curriculum.
Our experimental results are also bound to zero-shot settings, without exploring the effects of few-shot or test-time compute (Snell et al., 2024) prompting. Future work can investigate whether more sophisticated prompting enhances diagnostic accuracy. For example, prompting an LLM to explicitly infer a student’s thinking process before answering each check item may lead to more accurate judgments. We see our work as a first step toward automated cognitive skill assessment, and we invite the community to build upon these directions.
Acknowledgments
This work was supported by Algorithm LABS. This research (paper) used datasets from “The Open AI Dataset Project (AI-Hub, S. Korea).” All data information can be accessed through “AI-Hub (www.aihub.or.kr).”
Appendix A Appendix
Domains Cognitive Skills Description
Knowing Recall
Recall definitions, terminology, number properties, units of measurement,
geometric properties, and notation (e.g., $a\times b=ab,a+a+a=3a$).
Recognize
Recognize numbers, expressions, quantities, and shapes. Recognize entities
that are mathematically equivalent (e.g., equivalent familiar fractions,
decimals, and percents; different orientations of simple geometric figures).
Classify/Order Classify numbers, expressions, quantities, and shapes by common properties.
Compute
Carry out algorithmic procedures for +,,×,÷+,-$\times$,$\div$+ , - , × , ÷ or a combination of these
with whole numbers, fractions, decimals, and integers. Carry out
straightforward algebraic procedures.
Retrieve Retrieve information from graphs, tables, texts, or other sources.
Measure Use measuring instruments; and choose appropriate units of measurement.
Applying Determine
Determine efficient/appropriate operations, strategies, and tools for solving
problems for which there are commonly used methods of solution.
Represent/Model
Display data in tables or graphs; create equations, inequalities, geometric
figures, or diagrams that model problem situations; and generate equivalent
representations for a given mathematical entity or relationship.
Implement
Implement strategies and operations to solve problems involving familiar
mathematical concepts and procedures.
Reasoning Analyze
Determine, describe, or use relationships among numbers, expressions,
quantities, and shapes.
Integrate/Synthesize
Link different elements of knowledge, related representations, and
procedures to solve problems.
Evaluate Evaluate alternative problem solving strategies and solutions.
Draw conclusions Make valid inferences on the basis of information and evidence.
Generalize
Make statements that represent relationships in more general and more
widely applicable terms.
Justify Provide mathematical arguments to support a strategy or solution.
Table 3: Fifteen cognitive skills and their descriptions defined in the TIMSS 2019 framework.
Problem Types 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
% Agreement 95 70 96 95 89 68 74 86 80 80 100 88 40 57 78
Table 4: The inter-rater percentage absolute agreement of each problem type. The percentage indicates the ratio of unanimous verdicts in each teacher group.
A.1 Prompts
The blue text represents the programmatically filled arguments, and the orange text represents LLM-generated output.
A.2 System Prompt
# **Task Description**
You are a middle school math teacher tasked with evaluating students mathematical thinking skills based on their responses to math problems. Your goal is to analyze a given students response and determine whether they exhibit specific cognitive skills in solving the problem. Your evaluation must be **strict** and **evidence-based**, meaning that every verdict must be backed by direct evidence from the response. If no clear evidence exists, do not assume correctness.
## **Evaluation Categories**
For each thinking skill in the checklist, you must classify the students performance into one of the following categories:
- **Evident Yes**: The students response provides clear and explicit evidence that the check item is met. A direct quote from the response can confirm this.
- **Vague Yes**: The response suggests that the check item might be satisfied, but no specific part of the response directly proves it.
- **Evident No**: The response explicitly contradicts or fails to meet the check item, with clear evidence demonstrating the error or omission.
- **Vague No**: The response does not appear to satisfy the check item, but there is no direct evidence confirming whether the student considered it or not.
## **Input Format**
You will receive the following data:
- **Problem**: A math problem given to a student.
- **Answer**: The correct step-by-step solution.
- **Response**: The students response to the problem.
- **Check Items**: A set of specific skills to evaluate.
## **Output Format**
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
A.3 User Prompt
# **Task**
student responses to math problems, extract direct evidence, and strictly classify thinking skills according to the given categories.
Return a valid JSON object structured as follows:
‘‘‘json
"skills": [
"checkItem": "<Check Items [Label] and the Following Question>",
"evidence": "<Directly Quoted Part of Response>",
"explanation": "<Explanation About Why the Evidence Supports the Verdict>",
"verdict": "Evident Yes" | "Vague Yes" | "Evident No" | "Vague No"
**Problem:**
There is a two-digit natural number whose tens digit is $1$. If the number that changes the tens and ones digits of this natural number is $9$ less than $5$ times the first number, find the first number.
**Answer:**
If the number in the ones place is $x$, this natural number is $10+x$. The number where the tens digit and the ones digit are swapped is $110x+$ 1 because the tens digit is $x$ and the ones digit is $1$. The changed number is $9$ smaller than $5$ times the first number, so $910x+1=5(10+x)-$ 9 $4110x+1=5x+$ 41 $405x=$ 40 $x=8$ Therefore, the first number is $18$.
**Response:**
$1$ $x$ $x$ $1$ $910x+1=5(10+x)-$ 9 $910x+1=50+5x-$ 9 $41$ $405x=$ 40 $x=8$
**Check Items:**
- [recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?
- [recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?
- [compute_0] Has the student performed algebraic calculations correctly, such as expanding expressions and organizing similar terms?
- [determine_0] Has the student chosen the appropriate approach to solving the problem (e.g., establishing a linear equation)?
- [represent_0] Are two-digit numbers expressed correctly in the form 10x + y (or equivalent)?
- [implement_0] Does the student proceed consistently with the process of expanding the equation and moving terms to find x (or y) according to the planned solution strategy?
**Output:**
‘‘‘json
"skills": [
"checkItem": "[recognize_0] Is the student aware that the relationship between two numbers switched can be expressed as addition/subtraction/multiplication?",
"evidence": "10 x+1=5(10+x)-9",
"explanation": "The student sets up the equation 10x+1=5(10+x)-9, which shows they understand the relationship between the original and swapped numbers involves subtraction and multiplication.",
"verdict": "Evident Yes"
"checkItem": "[recognize_1] Has the student figured out all the elements necessary to solve the problem (original numbers, swapped numbers, difference or multiple relationships, etc.)?",
"evidence": "10 x+1=5(10+x)-9",
"explanation": "The student uses the correct equation to represent the relationship between the original and swapped numbers, indicating they have identified the necessary elements.",
"verdict": "Evident Yes"
Recall Recognize Classify/Order Compute Retrieve Determine Represent Implement
Evident Yes 507 506 92 584 9 441 154 319
Evident No 78 69 8 160 9 54 55 146
Vague Yes 58 54 2 48 0 25 6 29
Vague No 98 62 0 58 34 16 46 41
Student responses 582 583 51 796 52 482 261 535
Table 5: Distribution of verdicts and number of diagnosed student responses in each cognitive skill. Note that the number of student responses can be larger than the sum of the four labels because some diagnostic checklists have two items for the same cognitive skill.
Models Precision Recall F1 Accuracy Overconfidence Underconfidence
DeepSeek-R1 .447 .467 .442 .773 .901 .028
DeepSeek-V3 .406 .461 .417 .714 .710 .107
GPT-4o-img .427 .494 .448 .743 .738 .070
GPT-4o .398 .476 .412 .672 .703 .110
GPT-4o-Mini .345 .397 .352 .622 .791 .102
o1-Preview .338 .410 .354 .711 .757 .117
Llama-3.3-70B .416 .468 .400 .640 .534 .189
Llama-3.1-8B .283 .265 .257 .642 .695 .138
Llama-3.1-405B .415 .438 .384 .624 .451 .186
Claude-3.5-Sonnet-img .326 .365 .330 .690 .680 .140
Claude-3.5-Sonnet .329 .387 .332 .654 .676 .173
Gemini-1.5-Flash-img .431 .484 .429 .702 .593 .129
Gemini-1.5-Flash-8b .287 .341 .279 .558 .579 .196
Gemini-1.5-Flash .432 .490 .432 .679 .582 .155
Gemini-1.5-Pro .430 .496 .441 .707 .721 .132
Gemini-2.0-Flash-Thinking .455 .492 .421 .690 .589 .186
Mistral-7B .384 .369 .347 .714 .805 .112
Qwen-2.5-7B .278 .270 .271 .705 .842 .075
Qwen-2.5-72B .415 .459 .416 .710 .631 .115
Table 6: Performance comparison of all 16 LLMs tested on skill diagnosis tasks.
Figure 5: Illustrative examples of diagnosis check items and student responses that LLMs failed to diagnose correctly. Evidence for human judgment is marked with a red box.